OpenStudy (masonsurfs):

Can someone please help me find the solution to the systems of equations: 3x+2y+z=7 5x+5y+4z=3 3x+2y+3z=1 I'm looking for an explanation because I have been stuck on this question.

OpenStudy (seratul):

Okay, you can use elimination with the first and third equation. Subtract those 2 and tell me what you get.

OpenStudy (masonsurfs):

I got 2z=6

OpenStudy (seratul):

It should be -2z=6 Solve for z.

OpenStudy (masonsurfs):

Okay, z=-3

OpenStudy (seratul):

Okay, no rewrite all three of the equations but with -3 for z.

OpenStudy (masonsurfs):

3x+2y-3=7 5x+5y-12=3 3x+2y-9=1

OpenStudy (masonsurfs):

@Seratul

OpenStudy (seratul):

Simplify each equation.

OpenStudy (masonsurfs):

3x+2y=10 5x+5y=15 3x+2y=10

OpenStudy (seratul):

Okay, take the second equation and isolate the y. Make sure you divide by 5 at the end.

OpenStudy (masonsurfs):

y=-x+3

OpenStudy (seratul):

Wait, my bad, just divide by 5.

OpenStudy (masonsurfs):

Is that correct?

OpenStudy (seratul):

I meant to say just divide by 5 so it should be x+y=3 If you multiply by 2, you get 2x+2y=6 Now you can subtract 3x+2y=10 and 2x+2y=6. Tell me what you get. We are doing this so we can eliminate a variable.

OpenStudy (masonsurfs):

Okay so, x=4

OpenStudy (seratul):

Yep. Now plug in both x and z and find y.

OpenStudy (masonsurfs):

3(4)+2y-3=7 12+2y-3=7 9+2y=7 2y=-2 y=-1 ?

OpenStudy (seratul):

Yep. There you go.

OpenStudy (masonsurfs):

Great, thank you so much for the help!

OpenStudy (seratul):

np